# PICore: Physics-Informed Unsupervised Coreset Selection for Data Efficient Neural Operator Training

### v1

@@ 6.2.1 Unsupervised Coreset Selection
We compare PICore to three unsupervised coreset selection methods: k-means clustering, cosine similarity, and Herding (Chen et al., 2012). For k-means clustering we use $k=\beta N$ clusters, and choose the data points closest to those clusters. For cosine similarity, we evaluate the cosine similarity between all pairs of points, and perform greedy selection to choose the coreset. We report direct comparison of the test NRMSE for both methods in tables  5, 6, 7, and 8. The results show that PICore consistently matches or outperforms the unsupervised baselines across all tested coreset sizes (20% to 80%) and neural operator architectures (FNO and UNO). For instance, on the Advection dataset at 20% coreset size, PICore with the EL2N algorithm achieves a test NRMSE of $3.29\times 10^{-2}$, outperforming cosine similarity $3.39\times 10^{-2}$ and herding $3.46\times 10^{-2}$. Similar patterns are observed on the other datasets, indicating that PICore’s selection strategy generalizes well across both time-dependent and stationary PDEs compared to other unsupervised coreset selection strategies. We also note that these trends hold across neural operator architectures, with PICore outperforming unsupervised methods with both FNO and UNO architectures. While FNO does consistently outperform UNO across datasets (except Navier Stokes Incompressible), this is due to the architecture differences and not due to PICore (as shown by the increase in NRMSE for UNO on the non-coreset baseline).
@@
Operator Method Algorithm 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Unsupervised kmeans $3.85\times 10^{-2}$ $3.48\times 10^{-2}$ $2.98\times 10^{-2}$ $2.59\times 10^{-2}$ $2.35\times 10^{-2}$ $2.13\times 10^{-2}$
cosine $3.39\times 10^{-2}$ $3.17\times 10^{-2}$ $2.76\times 10^{-2}$ $2.52\times 10^{-2}$ $2.35\times 10^{-2}$ $2.13\times 10^{-2}$
herding $3.46\times 10^{-2}$ $3.14\times 10^{-2}$ $\mathbf{2.68\times 10^{-2}}$ $2.46\times 10^{-2}$ $2.25\times 10^{-2}$ $2.13\times 10^{-2}$
PICore craig $3.66\times 10^{-2}$ $3.28\times 10^{-2}$ $2.94\times 10^{-2}$ $2.56\times 10^{-2}$ $2.32\times 10^{-2}$ $2.13\times 10^{-2}$
gradmatch $3.38\times 10^{-2}$ $3.22\times 10^{-2}$ $2.70\times 10^{-2}$ $2.44\times 10^{-2}$ $\mathbf{2.20\times 10^{-2}}$ $2.13\times 10^{-2}$
adacore $5.08\times 10^{-2}$ $4.00\times 10^{-2}$ $3.26\times 10^{-2}$ $2.70\times 10^{-2}$ $2.40\times 10^{-2}$ $2.13\times 10^{-2}$
el2n $\mathbf{3.29\times 10^{-2}}$ $\mathbf{3.03\times 10^{-2}}$ $2.71\times 10^{-2}$ $2.49\times 10^{-2}$ $2.35\times 10^{-2}$ $2.13\times 10^{-2}$
graNd $3.46\times 10^{-2}$ $3.04\times 10^{-2}$ $2.69\times 10^{-2}$ $\mathbf{2.40\times 10^{-2}}$ $2.25\times 10^{-2}$ $2.13\times 10^{-2}$
UNO Unsupervised kmeans $1.61\times 10^{-1}$ $1.52\times 10^{-1}$ $1.44\times 10^{-1}$ $1.39\times 10^{-1}$ $1.13\times 10^{-1}$ $7.21\times 10^{-2}$
cosine $1.53\times 10^{-1}$ $1.46\times 10^{-1}$ $1.43\times 10^{-1}$ $1.32\times 10^{-1}$ $9.21\times 10^{-2}$ $7.21\times 10^{-2}$
herding $1.54\times 10^{-1}$ $1.46\times 10^{-1}$ $1.42\times 10^{-1}$ $1.27\times 10^{-1}$ $9.51\times 10^{-2}$ $7.21\times 10^{-2}$
PICore craig $1.60\times 10^{-1}$ $1.48\times 10^{-1}$ $1.44\times 10^{-1}$ $1.29\times 10^{-1}$ $9.67\times 10^{-2}$ $7.21\times 10^{-2}$
gradmatch $1.55\times 10^{-1}$ $1.47\times 10^{-1}$ $1.43\times 10^{-1}$ $\mathbf{1.26\times 10^{-1}}$ $9.06\times 10^{-2}$ $7.21\times 10^{-2}$
adacore $2.25\times 10^{-1}$ $1.71\times 10^{-1}$ $1.47\times 10^{-1}$ $1.31\times 10^{-1}$ $1.24\times 10^{-1}$ $7.21\times 10^{-2}$
el2n $\mathbf{1.52\times 10^{-1}}$ $\mathbf{1.45\times 10^{-1}}$ $1.42\times 10^{-1}$ $1.30\times 10^{-1}$ $9.89\times 10^{-2}$ $7.21\times 10^{-2}$
graNd $1.54\times 10^{-1}$ $1.46\times 10^{-1}$ $\mathbf{1.42\times 10^{-1}}$ $1.33\times 10^{-1}$ $\mathbf{9.04\times 10^{-2}}$ $7.21\times 10^{-2}$
Table 5: Test NRMSE on the Advection dataset at resolution 64 across varying coreset percentages (20%–100%) between unsupervised and PICore-based coreset selection methods using both FNO and UNO architectures.
Operator Method Algorithm 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Unsupervised kmeans $1.98\times 10^{-2}$ $1.36\times 10^{-2}$ $1.02\times 10^{-2}$ $6.35\times 10^{-3}$ $4.45\times 10^{-3}$ $3.61\times 10^{-3}$
cosine $1.88\times 10^{-2}$ $1.31\times 10^{-2}$ $9.90\times 10^{-3}$ $5.66\times 10^{-3}$ $4.31\times 10^{-3}$ $3.61\times 10^{-3}$
herding $1.83\times 10^{-2}$ $1.35\times 10^{-2}$ $8.27\times 10^{-3}$ $5.66\times 10^{-3}$ $4.17\times 10^{-3}$ $3.61\times 10^{-3}$
PICore craig $1.93\times 10^{-2}$ $1.56\times 10^{-2}$ $1.10\times 10^{-2}$ $6.68\times 10^{-3}$ $4.40\times 10^{-3}$ $3.61\times 10^{-3}$
gradmatch $1.91\times 10^{-2}$ $1.15\times 10^{-2}$ $8.08\times 10^{-3}$ $5.61\times 10^{-3}$ $4.24\times 10^{-3}$ $3.61\times 10^{-3}$
adacore $2.86\times 10^{-2}$ $2.32\times 10^{-2}$ $1.46\times 10^{-2}$ $8.07\times 10^{-3}$ $4.90\times 10^{-3}$ $3.61\times 10^{-3}$
el2n $\mathbf{1.81\times 10^{-2}}$ $\mathbf{1.12\times 10^{-2}}$ $\mathbf{8.07\times 10^{-3}}$ $\mathbf{5.49\times 10^{-3}}$ $\mathbf{4.07\times 10^{-3}}$ $3.61\times 10^{-3}$
graNd $1.87\times 10^{-2}$ $1.26\times 10^{-2}$ $8.15\times 10^{-3}$ $5.50\times 10^{-3}$ $4.08\times 10^{-3}$ $3.61\times 10^{-3}$
UNO Unsupervised kmeans $3.12\times 10^{-2}$ $2.70\times 10^{-2}$ $2.20\times 10^{-2}$ $1.81\times 10^{-2}$ $1.54\times 10^{-2}$ $1.68\times 10^{-2}$
cosine $2.95\times 10^{-2}$ $2.39\times 10^{-2}$ $2.10\times 10^{-2}$ $1.76\times 10^{-2}$ $1.54\times 10^{-2}$ $1.68\times 10^{-2}$
herding $2.96\times 10^{-2}$ $2.59\times 10^{-2}$ $2.11\times 10^{-2}$ $1.72\times 10^{-2}$ $1.52\times 10^{-2}$ $1.68\times 10^{-2}$
PICore craig $3.28\times 10^{-2}$ $3.01\times 10^{-2}$ $2.33\times 10^{-2}$ $1.90\times 10^{-2}$ $1.57\times 10^{-2}$ $1.68\times 10^{-2}$
gradmatch $2.96\times 10^{-2}$ $2.43\times 10^{-2}$ $2.08\times 10^{-2}$ $1.75\times 10^{-2}$ $1.56\times 10^{-2}$ $1.68\times 10^{-2}$
adacore $5.09\times 10^{-2}$ $3.96\times 10^{-2}$ $3.17\times 10^{-2}$ $2.21\times 10^{-2}$ $1.72\times 10^{-2}$ $1.68\times 10^{-2}$
el2n $2.95\times 10^{-2}$ $2.47\times 10^{-2}$ $2.10\times 10^{-2}$ $1.73\times 10^{-2}$ $\mathbf{1.51\times 10^{-2}}$ $1.68\times 10^{-2}$
graNd $\mathbf{2.84\times 10^{-2}}$ $\mathbf{2.36\times 10^{-2}}$ $\mathbf{2.06\times 10^{-2}}$ $\mathbf{1.72\times 10^{-2}}$ $1.57\times 10^{-2}$ $1.68\times 10^{-2}$
Table 6: Test NRMSE on the Burgers dataset at resolution 64 across varying coreset percentages (20%–100%) between unsupervised and PICore-based coreset selection methods using both FNO and UNO architectures.
Operator Method Algorithm 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Unsupervised kmeans $1.28\times 10^{-1}$ $1.18\times 10^{-1}$ $9.62\times 10^{-2}$ $7.83\times 10^{-2}$ $6.78\times 10^{-2}$ $5.96\times 10^{-2}$
cosine $1.29\times 10^{-1}$ $1.15\times 10^{-1}$ $9.88\times 10^{-2}$ $7.78\times 10^{-2}$ $6.85\times 10^{-2}$ $5.96\times 10^{-2}$
herding $1.29\times 10^{-1}$ $1.12\times 10^{-1}$ $\mathbf{9.35\times 10^{-2}}$ $7.82\times 10^{-2}$ $6.80\times 10^{-2}$ $5.96\times 10^{-2}$
PICore craig $1.36\times 10^{-1}$ $1.20\times 10^{-1}$ $1.05\times 10^{-1}$ $8.20\times 10^{-2}$ $6.85\times 10^{-2}$ $5.96\times 10^{-2}$
gradmatch $1.33\times 10^{-1}$ $1.13\times 10^{-1}$ $9.97\times 10^{-2}$ $7.92\times 10^{-2}$ $\mathbf{6.68\times 10^{-2}}$ $5.96\times 10^{-2}$
adacore $1.66\times 10^{-1}$ $1.33\times 10^{-1}$ $1.11\times 10^{-1}$ $8.00\times 10^{-2}$ $6.88\times 10^{-2}$ $5.96\times 10^{-2}$
el2n $\mathbf{1.25\times 10^{-1}}$ $1.12\times 10^{-1}$ $9.44\times 10^{-2}$ $\mathbf{7.77\times 10^{-2}}$ $6.84\times 10^{-2}$ $5.96\times 10^{-2}$
graNd $1.32\times 10^{-1}$ $\mathbf{1.11\times 10^{-1}}$ $9.75\times 10^{-2}$ $7.82\times 10^{-2}$ $6.81\times 10^{-2}$ $5.96\times 10^{-2}$
UNO Unsupervised kmeans $1.45\times 10^{-1}$ $1.22\times 10^{-1}$ $1.06\times 10^{-1}$ $9.21\times 10^{-2}$ $7.77\times 10^{-2}$ $6.71\times 10^{-2}$
cosine $1.31\times 10^{-1}$ $1.14\times 10^{-1}$ $1.00\times 10^{-1}$ $8.61\times 10^{-2}$ $7.74\times 10^{-2}$ $6.71\times 10^{-2}$
herding $1.34\times 10^{-1}$ $1.17\times 10^{-1}$ $1.01\times 10^{-1}$ $8.81\times 10^{-2}$ $7.80\times 10^{-2}$ $6.71\times 10^{-2}$
PICore craig $1.49\times 10^{-1}$ $1.32\times 10^{-1}$ $1.14\times 10^{-1}$ $9.31\times 10^{-2}$ $8.00\times 10^{-2}$ $6.71\times 10^{-2}$
gradmatch $1.29\times 10^{-1}$ $\mathbf{1.10\times 10^{-1}}$ $9.87\times 10^{-2}$ $8.44\times 10^{-2}$ $7.69\times 10^{-2}$ $6.71\times 10^{-2}$
adacore $1.84\times 10^{-1}$ $1.51\times 10^{-1}$ $1.25\times 10^{-1}$ $9.86\times 10^{-2}$ $8.21\times 10^{-2}$ $6.71\times 10^{-2}$
el2n $1.33\times 10^{-1}$ $1.14\times 10^{-1}$ $9.83\times 10^{-2}$ $8.45\times 10^{-2}$ $\mathbf{7.55\times 10^{-2}}$ $6.71\times 10^{-2}$
graNd $\mathbf{1.28\times 10^{-1}}$ $1.12\times 10^{-1}$ $\mathbf{9.67\times 10^{-2}}$ $\mathbf{8.42\times 10^{-2}}$ $7.61\times 10^{-2}$ $6.71\times 10^{-2}$
Table 7: Test NRMSE on the Darcy dataset at resolution 64 across varying coreset percentages (20%–100%) between unsupervised and PICore-based coreset selection methods using both FNO and UNO architectures.
Operator Method Algorithm 20.0% 30.0% 40.0% 60.0% 80.0% 100.0%
FNO Unsupervised kmeans $\mathbf{8.50\times 10^{-2}}$ $1.86\times 10^{-2}$ $1.37\times 10^{-2}$ $8.76\times 10^{-3}$ $6.56\times 10^{-3}$ $5.56\times 10^{-2}$
cosine $1.37\times 10^{-1}$ $2.30\times 10^{-2}$ $1.32\times 10^{-2}$ $8.60\times 10^{-3}$ $6.51\times 10^{-3}$ $5.56\times 10^{-2}$
herding $1.69\times 10^{-1}$ $1.81\times 10^{-2}$ $1.23\times 10^{-2}$ $8.85\times 10^{-3}$ $6.59\times 10^{-3}$ $5.56\times 10^{-2}$
PICore craig $1.36\times 10^{-1}$ $3.21\times 10^{-2}$ $1.50\times 10^{-2}$ $8.88\times 10^{-3}$ $6.71\times 10^{-3}$ $5.56\times 10^{-2}$
gradmatch $1.26\times 10^{-1}$ $1.80\times 10^{-2}$ $1.24\times 10^{-2}$ $8.06\times 10^{-3}$ $6.44\times 10^{-3}$ $5.56\times 10^{-2}$
adacore $2.72\times 10^{-1}$ $7.04\times 10^{-2}$ $1.68\times 10^{-2}$ $9.79\times 10^{-3}$ $6.96\times 10^{-3}$ $5.56\times 10^{-2}$
el2n $1.19\times 10^{-1}$ $\mathbf{1.67\times 10^{-2}}$ $\mathbf{1.19\times 10^{-2}}$ $\mathbf{7.94\times 10^{-3}}$ $6.36\times 10^{-3}$ $5.56\times 10^{-2}$
graNd $1.12\times 10^{-1}$ $1.81\times 10^{-2}$ $1.23\times 10^{-2}$ $8.00\times 10^{-3}$ $\mathbf{6.34\times 10^{-3}}$ $5.56\times 10^{-2}$
UNO Unsupervised kmeans $2.86\times 10^{-2}$ $2.40\times 10^{-2}$ $2.09\times 10^{-2}$ $1.67\times 10^{-2}$ $1.42\times 10^{-2}$ $1.23\times 10^{-2}$
cosine $2.64\times 10^{-2}$ $2.21\times 10^{-2}$ $1.93\times 10^{-2}$ $1.60\times 10^{-2}$ $1.40\times 10^{-2}$ $1.23\times 10^{-2}$
herding $2.65\times 10^{-2}$ $2.21\times 10^{-2}$ $1.94\times 10^{-2}$ $1.60\times 10^{-2}$ $1.39\times 10^{-2}$ $1.23\times 10^{-2}$
PICore craig $2.85\times 10^{-2}$ $2.42\times 10^{-2}$ $2.14\times 10^{-2}$ $1.69\times 10^{-2}$ $1.44\times 10^{-2}$ $1.23\times 10^{-2}$
gradmatch $\mathbf{2.59\times 10^{-2}}$ $\mathbf{2.20\times 10^{-2}}$ $\mathbf{1.92\times 10^{-2}}$ $1.60\times 10^{-2}$ $1.40\times 10^{-2}$ $1.23\times 10^{-2}$
adacore $3.79\times 10^{-2}$ $3.07\times 10^{-2}$ $2.57\times 10^{-2}$ $1.90\times 10^{-2}$ $1.51\times 10^{-2}$ $1.23\times 10^{-2}$
el2n $2.60\times 10^{-2}$ $2.20\times 10^{-2}$ $1.93\times 10^{-2}$ $1.59\times 10^{-2}$ $1.39\times 10^{-2}$ $1.23\times 10^{-2}$
graNd $2.60\times 10^{-2}$ $2.21\times 10^{-2}$ $1.93\times 10^{-2}$ $\mathbf{1.59\times 10^{-2}}$ $\mathbf{1.39\times 10^{-2}}$ $1.23\times 10^{-2}$
Table 8: Test NRMSE on the Navier Stokes Incompressible dataset at resolution 64 across varying coreset percentages (20%–100%) between unsupervised and PICore-based coreset selection methods using both FNO and UNO architectures.
6.2.2 Comparison between Supervised Coreset Selection and PICore
To better understand the differences between supervised coreset selection and PICore, we analyze how well each method covers the input space by computing the average distance from coreset points to their centroid, which serves as a proxy for spread or diversity. As shown in Figure 2, this distance is nearly identical across datasets and neural operators (FNO and UNO), with overlapping standard error bars with differences decreasing as the PDE complexity increases (Advection to Navier Stokes). This suggests that PICore selects coresets that are as well-distributed as those from supervised methods, despite not using labeled data. The comparable coverage indicates that differences in downstream performance likely arise from the type of points selected rather than their spatial distribution.
(a) Advection
(b) Burgers
(c) Darcy
(d) Navier Stokes
Figure 2: Average centroid distances across datasets for FNO and UNO.
@@ 7 Limitations and Future Work
One limitation of our work lies in its reliance on existing coreset selection algorithms. Methods such as CRAIG and AdaCore were originally developed under convexity assumptions and gradient and Hessian approximations and have shown strong performance in domains like image classification. However, applying such methods to complex PDE datasets, where the loss landscape is often highly non-convex, is less well understood. As a result, PICore’s performance at low selection ratios may be constrained by the suboptimal behavior of these algorithms. For example, using the Hutchinson Hessian approximation for Adacore on only the last layer of the neural operator consistently results in poorer accuracies on the selected coresets than other coreset selection algorithms. Thus, practitioners are advised to use GradMatch, EL2N, or graNd as they operate under fewer model dependent assumptions and approximations. Future work could improve PICore by designing coreset selection algorithms tailored specifically for neural operator learning, by potentially leveraging inductive biases in operator networks. Additionally, while PICore can work with any neural operator architecture and input, we only use PICore on one input resolution with uniform geometries. Integrating multi-resolution data with arbitrary geometries (and adjusting the neural operator to handle them) could improve its generalization capability while maintaining efficiency.
@@ Appendix A PDE Datasets
For our experiments, we use several differential equation training sets to evaluate our algorithm. Each of these is used at an input grid resolution of 64. For the Advection, Burgers, and Darcy Flow equations, we generate datasets using code provided by Takamoto et al. (2022). For the Navier-Stokes Incompressible equation dataset, we generate data from Li et al. (2020). All data generation occurs at the largest resolution, and we downsample by selecting equally spaced points to obtain lower resolution solutions.
@@ A.1 Advection
where each $n_{i}$ is drawn uniformly from the range of integers from 1 to 8, $N$ is the number of waves, and the amplitudes $A_{i}\in[0,1]$ and phases $\phi_{i}\in(0,2\pi)$ are chosen at random. After assembly of $u_{0}(x)$, we apply with 10% probability each a pointwise absolute-value operation or multiplication by a smooth window function.
@@ A.2 Burger’s Equation
where $u_{L}$ is a characteristic velocity scale. In analogy with the Navier–Stokes equations, $R>1$ indicates a regime dominated by nonlinear steepening and potential shock formation, whereas $R<1$ corresponds to diffusion-dominated smooth dynamics.

### v2

@@ 6.2.1 Unsupervised Coreset Selection
We compare PICore to three unsupervised coreset selection methods: k-means clustering, cosine similarity, and Herding (Chen et al., 2012). For k-means clustering we use $k=\beta N$ clusters, and choose the data points closest to those clusters. For cosine similarity, we evaluate the cosine similarity between all pairs of points, and perform greedy selection to choose the coreset. We report direct comparison of the test NRMSE for both methods in Figures  4, 5, 6, and 7 in Section C.1. The results show that PICore consistently matches or outperforms the unsupervised baselines across all tested coreset sizes (20% to 80%) and neural operator architectures (FNO and UNO). For instance, on the Advection dataset at 20% coreset size, PICore with the EL2N algorithm achieves a test NRMSE of $3.29\times 10^{-2}$, outperforming cosine similarity $3.39\times 10^{-2}$ and herding $3.46\times 10^{-2}$. Similar patterns are observed on the other datasets, indicating that PICore’s selection strategy generalizes well across both time-dependent and stationary PDEs compared to other unsupervised coreset selection strategies. We also note that these trends hold across neural operator architectures, with PICore outperforming unsupervised methods with both FNO and UNO architectures. While FNO does consistently outperform UNO across datasets (except Navier Stokes Incompressible), this is due to the architecture differences and not due to PICore, as shown by the increase in NRMSE for UNO on the non-coreset baseline.
@@
6.2.2 Convergence of Coreset Selection vs Active Learning
A key distinction between coreset selection and active learning lies in their approach to data selection, which in turn affects their convergence speed. This iterative nature can be suboptimal, as it can lead to selecting redundant data points (Li et al., 2024a). While some works have shown superior convergence of active learning methods (Haimovich et al., 2024), these are under specific optimizer settings and in easier image classification domains.
[…]
Figure 2: Training convergence of coreset selection v.s. active learning using FNO at a 20% selection ratio.
Figure 3: Training convergence of coreset selection v.s. active learning using UNO at a 20% selection ratio.
@@ 7 Limitations and Future Work
One key limitation of PICore is the physics-informed loss requires that the underlying PDE is known and differentiable, which may not always be possible. However, practitioners can circumvent this by estimating the PDE analytically by using a trustworthy auxiliary model via domain knowledge to construct the PDE and using subdifferentials to estimate its derivatives, or numerically by employing data-driven surrogates or weak-form formulations that approximate the governing equations and their derivatives from observations. If such methods are not possible, alternative unsupervised criteria directly from observation, such as the distribution of the initial conditions, can be used with lower accuracy. Another limitation of our work lies in its reliance on existing coreset selection algorithms. Methods such as CRAIG and AdaCore were originally developed under convexity assumptions and gradient and Hessian approximations and have shown strong performance in domains like image classification. However, applying such methods to complex PDE datasets, where the loss landscape is often highly non-convex, is less well understood. As a result, PICore’s performance at low selection ratios may be constrained by the suboptimal behavior of these algorithms. For example, using the Hutchinson Hessian approximation for AdaCore on only the last layer of the neural operator consistently results in poorer accuracies on the selected coresets than other coreset selection algorithms. Thus, practitioners are advised to use GradMatch, EL2N, or graNd as they operate under fewer model dependent assumptions and approximations. Future work could improve PICore by designing coreset selection algorithms tailored specifically for neural operator learning, by potentially leveraging inductive biases in operator networks. Additionally, while PICore can work with any neural operator architecture and input, we only use PICore on one input resolution with uniform geometries. Integrating multi-resolution data with arbitrary geometries could improve its generalization capability while maintaining efficiency.
@@ Appendix A PDE Datasets
For our experiments, we use several differential equation training sets to evaluate our algorithm. Each of these is used at an input grid resolution of 64. For the Advection, Burgers, and Darcy Flow equations, we generate datasets using code provided by Takamoto et al. (2022). For the Navier-Stokes Incompressible equation dataset, we generate data from Li et al. (2020).
@@ A.1 Advection
where each $n_{i}$ is drawn uniformly from the range of integers from 1 to 8, $N$ is the number of waves, and the amplitudes $A_{i}\in[0,1]$ and phases $\phi_{i}\in(0,2\pi)$ are chosen at random. After assembly of $u_{0}(x)$, we apply with 10% probability each a pointwise absolute‐value operation or multiplication by a smooth window function.
@@ A.2 Burger’s Equation
where $u_{L}$ is a characteristic velocity scale. In analogy with the Navier–Stokes equations, $R>1$ indicates a regime dominated by nonlinear steepening and potential shock formation, whereas $R<1$ corresponds to diffusion‐dominated smooth dynamics.